User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair/Presenting: Bryn

In Attendance: Liz, Nick, Tiege, Lucas, Ross, Jonathan, Jesse, Jamie, Ellen, Jim, Lucia

Apologies: Arwin




General Issues

  • LBA this weekend - keep skype open and check the LBA chat, let atnf people know when you come on shift
    • LBA wiki has page to note issues
  • Check disk space before starting the experiment
    • Someone is always at Yg - if unsure, can check.
    • Data Volume for 256 Mb LBA is 3TB/Day - use this as rule of thumb if no .sum file
  • Watchdog alarms on opsII - has anyone noticed any alarms not sounding?
  • Identity on opsII - still need to send emails from ops2
    • Now named ops8
    • ssh into ops2 to send e-mails

Issues from Handover notes

* Yg VPN still dodgy
  * Someone notify Brett. 
* AUG031 schedule issue.
  * Caused by renaming issues with IVS
* AUG031 ke and yg fs lockup - required reboots
* HOB006 Ho tsys overflowing again - is there an issue with the CAL causing it to drop out mid experiments?
  * Happens several time through the experiment - could be an issue with one of Jim's scripts firing the cals.
  * Check with Jim P. (Desktop 3 above receiver selection, check with AT queue to see if a command is queued).
* Hb not coming out of wind stows properly - caused by either long wind stows or immediate wind stow after release
  * If drives trip off while trying to stow (doesn't reach stow position), the fs can get stuck in a stow loop. Needed to restart the fs.
  * Multiple onsource slewing errors could be programmed to sound a major alarm.
* R4762 ke ALARM: error s5 -104 rfpcn: time-out, connection closed occurred a lot during this experiment
  * Resetting rxmon and monica several times may have fixed this error.
* t2113 ke MET sensor disconnected - GA upgrading GNSS receiver & new interface is required
  * Seems to be working now - check to make sure met data makes sense (e.g. is changing with time)
* Yg and Ke webcams faulty/no connections
  * Yg antenna camera is also broken.
  * Working as of this meeting - may be intermittent
* Dynamic observing @ ke: ALARM: error m5 -900 not while recording or playing
  * Changed to a beep
  * Possibly caused by autocorrelation data being collected during recording **(Jamie please check)**
  • “ ERROR ch -4 Device v3 timed-out on response from MATCN ” means that something is wrong with the mk4 VCs and the mk4 rack (white one with 'VC' lights) needs to be restarted manually at Mt P. This error appeared after setupsx when changing the 26m back to the mk4 from the mk5/dbbc. The rack can be reset by turning off, waiting then turning on again and waiting (red lights everywhere until you setupsx again). The on/off switch is at the top of the rack. If this doesn't fix the matcn error then you needs to call Brett.
    • You can run up the setupsx to check the mk4 while Jim P. is observing vela.
/home/www/auscope/opswiki/data/pages/operations/agenda2016_10_27.txt · Last modified: 2016/10/27 02:40 by Ellen Manning